OpenStudy (anonymous):

simplify \[\sqrt{81}\div \sqrt{7}\] simplify \[\sqrt{81}\div \sqrt{7}\] @Mathematics

OpenStudy (anonymous):

9/\[\sqrt{7}\]

OpenStudy (anonymous):

\[\frac{9}{\sqrt{7}}=\frac{9\sqrt{7}}{{7}}\]